# Global dynamical structures from infinitesimal data

### v1

@@ Discussion
We have proposed VERT, a framework for using observed trajectory data to estimate the attracting sets that govern their asymptotic behavior (Fig. 1). We show that the relative fiberwise proximity of a measured state to the locally strongest attracting set can be estimated using a model agnostic, infinitesimally (i.e., pointwise) motivated linear calculation, applied to only a spatially local subsample of the full dataset, enabling inference of the global qualitative behavior of the system without recourse to a particular global model of the dynamics. The VERT framework complements existing approaches for learning predictive dynamics models from data, which typically assume post-transient training data [14], by providing a procedure to isolate transient trajectory segments from the full dataset. Because VERT does not require fitting a global model, this can be done in pre-processing. VERT further affords the ability to discover dynamical structure (the locus of attracting sets) from data that may be less amenable to global autoregressive model fitting due to the complexity of the underlying dynamics (e.g., presence of multiscale and discontinuous phenomena) and obstacles to model interpretation (e.g., systems that lack ground truth constitutive models). Stable attracting sets provide critical insight into the mechanisms underlying the dynamics of biological systems including locomotion [42], gene regulatory networks [48], and neural population dynamics [49]. We thus expect broad utility of VERT for learning models of biological and engineering systems, where the behavior of observed trajectories is strongly influenced by persistent disturbances and internal closed loop control processes.
@@
We demonstrated the efficacy of VERT for accurately identifying attracting sets in synthetic data from dynamical systems with both parallel (Fig. 3) and sequential (Fig. 4) hierarchies, and experimentally measured human running kinematics (Fig. 5). Our results show that when subsystems with different characteristic time scales are simultaneously active, VERT estimates the distance to the attracting set of the locally dominant subsystem, affording sensitivity to and potentially inference of such parallel hierarchies. In systems with hybrid or jump-like behaviors, VERT detects the locally stable attracting sets on continuous subdomains of the global attracting set. In the human running data, which exhibits features of both parallel and sequential dynamical hierarchy, this sensitivity suggests inference of potential control modules during the stance phase of running. The sublevel sets of the distance estimator accurately predict the phases of stance where ground truth kinematic observables exhibit template behavior, e.g., SLIP-like leg length dynamics and vertical GRF profile. Mounting experimental evidence suggests sensitivity of muscle spindle primary afferents to the rate of change of loading force as a mechanism for reflexive stabilization of postural perturbations [50]. The VERT predictions of post-transient template phases of stance are characterized by linear or constant accelerations of the ’template’ kinematic variables for the respective arm-torso and leg control modules, suggesting that the transient behavior may result from reflexive stabilization following impact. These observations support the possibility of using VERT to accurately predict muscle activation dynamics from motion capture data, a prospect with potential breakthrough applications for biomechanics, robotics, and simulation of human dynamics. VERT can further be used in experimental design to develop perturbation assays that untangle the roles of different postural control modules for stabilizing locomotion tasks.
@@
Integration of the VERT pipeline with tools for geometric and topological inference is a natural next step — discovered posture submanifolds likely (e.g., almost always in robotics) have dimensions higher than $1$, and are almost always highly nonlinear. This motivates developing a funnel to the computational pipes of Topological Data Analysis [51], allowing one to recognize, in a model-agnostic way, the topological structure of the underlying attracting set. As the attracting sets are still expected to be of relatively low dimensions, we conjecture that the available topological inferences (primarily, ranks of the homology groups) would be sufficient to achieve this goal. Once the underlying attracting set is recovered, the estimation of the template behavior — the dynamics on the attractor — could be achieved by the standard approximation tools [52, 53]. The promising results from the A-HM and A-vdP models motivate further work on inferences for systems with multiscale behavior. We showed that applying VERT to hypothesis-informed projections of the human postural state can provide insight into the dynamics of particular appendages and body segments. Learning such projections directly from the outputs of VERT is an exciting prospect. To ensure our analysis of the human running data was refined to the stance phase, we used the independently measured vertical GRF signal to segment the full trial into phases when the right leg was in contact. In principle, integrating detection of hybrid reset events into VERT should be possible. The saltation matrix characterizes the first order variational update at a reset [54], and can be estimated from the local sample covariance matrices, which are already computed and stored by VERT to estimate the vielbein. Future work will explore the prospect of using such estimates to localize guard surfaces in state space.
@@
Trajectories for the A-HM model were produced by integrating the model defined by Eqns. 1 & 2 using MATLAB’s ode45 function for the integration interval $[0:0.05:10]$. 10 polar coordinate pairs were sampled from the continuous uniform distributions $\rho\sim\mathcal{U}(2,4)$ and $\theta\sim\mathcal{U}(0,2\pi)$. For each polar coordinate pair $(\rho,\theta)$, five initial conditions were computed by sampling: $(\rho,\theta)+\mathcal{N}(0,0.3)$, producing a dataset of 50 trajectories. Trajectories were transformed to Cartesian coordinates before passing to VERT. To produce the plots in Fig. 3, control parameter combinations $(\gamma_{\rho},\gamma_{\delta})$ were produced from a uniformly spaced $50\times 50$ grid with $\gamma_{\rho},\gamma_{\delta}\in[1,3]$. VERT was used to compute the distance estimator for 10 trajectories. The MSE residual was computed over these 10 trajectories. For each control parameter combination, we performed 5 trials each with independently sampled initial conditions, and the MSE residuals for each trial were averaged to produce Figs. 3C,D. Details for computing the MSE residuals are described in the Supplement. Code and parameters to reproduce the plots in Fig. 3 are available in the Supplement.
@@
Trajectories for the A-VDP model were produced by integrating the model defined by Eqn. 3 using MATLAB’s ode45 function for the integration interval $[0:0.01:20]$. To produce the plots in Fig. 4, 20 initial conditions were independently sampled from the normal distributions: $x_{1}\sim\mathcal{N}(0.5,0.2)$, $x_{2}\sim\mathcal{N}(0.3,0.2)$, $x_{3}\sim\mathcal{N}(0.5,0.2)$, $x_{4}\sim\mathcal{N}(0.1,0.2)$. The numerical derivative $\dot{x}_{2}$ plotted in Fig. 4 was obtained using first order finite differences.
@@
The human running kinematics dataset consisted of a single continuous 4 min. trial (S3R1) obtained from the Dryad repository [45]. To parameterize the postural state of the right side of the body, we used the Euclidean distances between all pairwise combinations of the markers on the right side of the body as the position coordinates (see Supplement for marker labels), and the numerical derivatives of the position trajectories (obtained using finite differences) as the velocity coordinates, resulting in a state dimension of 2×($1822\times\binom{14}{2}=$ 182. The postural state trajectories were passed directly to the VERT pipeline. To produce the plots in Fig. 5, the vertical ground reaction force data was used to estimate right foot touchdown and liftoff events. The data was then aligned such that touchdown events occurred at $t=0s$. Mean trajectories for all variables over the stance phases were computed by linearly resampling the data on a uniform interval and computing the means across trials at each time step. The projected datasets were obtained following the above procedure for the marker sets used in the upper and lower body projections.
@@ Acknowledgements.
I Appendix
Proof of Theorem 1
We start with observing that $B(x_{*},0)=0$, and that, by the standard bound via derivative,
$|B(x,y)|=|B(x,y)-B(x_{*},0)|\leq C_{G}(|y|+|x-x_{*}|),$
so that $|B(x_{*},y_{*})|\leq C_{G}|y_{*}|$.
Expanding the definition of the estimator $f(x_{*},z_{*})=|H(x_{*},y_{*})v(x_{*},y_{*})|$, we obtain (abbreviating $A_{1}(x_{*})$ to $A_{1}$
$Hv=\left(\begin{pmatrix}0&0\\ 0&A_{1}\end{pmatrix}+G(x_{*},y_{*})\right)\left(\begin{pmatrix}u_{0}\\ A_{1}y_{*}\end{pmatrix}+B(x_{*},y_{*})\right).$
Opening the brackets, we obtain the following sum:
$\begin{array}[]{l}\begin{pmatrix}0\\ A_{1}^{2}y_{*}\end{pmatrix}+\begin{pmatrix}0&0\\ 0&A_{1}\end{pmatrix}B(x_{*},y_{*})+G(x_{*},y_{*})\begin{pmatrix}u_{0}\\ 0\end{pmatrix}+\\ G(x_{*},y_{*})\begin{pmatrix}0\\ A_{1}y_{*}\end{pmatrix}+G(x_{*},y_{*})B(x_{*},y_{*}).\end{array}$
The norm of the second term is bounded by $C_{+}C_{G}|y_{*}|$ (recall that $|B|\leq C_{G}|y_{*}|$); the norm of the third term is bounded by $C_{G}v_{0}$; of the fourth, $C_{G}C_{+}|y_{*}|$, and of the fifth, $C_{G}^{2}|y_{*}|$.
Further, as
$C_{-}^{2}|y_{*}|\leq\left|A_{1}^{2}y_{*}\right|\leq C_{+}^{2}|y_{*}|,$
we obtain our result after a straightforward algebraic manipulation. $\square$

### v2

@@ Discussion
We have proposed VERT, a framework for using observed trajectory data to estimate the attracting sets (termed ”postures”) that govern their asymptotic behavior (the restriction dynamics termed a ”template” Fig. 1). We show that the relative fiberwise proximity of a measured state to the locally strongest attracting set can be estimated using a model agnostic, infinitesimally (i.e., pointwise) motivated linear calculation, applied to only a spatially local subsample of the full dataset, enabling inference of the global qualitative behavior of the system without recourse to a particular global model of the dynamics. The VERT framework complements existing approaches for learning predictive dynamics models from data, which typically assume post-transient training data [42], by providing a procedure to isolate transient trajectory segments from the full dataset. Because VERT does not require fitting a global model, this can be done in pre-processing. VERT further affords the ability to discover dynamical structure (the locus of attracting sets) from data that may be less amenable to global autoregressive model fitting due to the complexity of the underlying dynamics (e.g., presence of multiscale and discontinuous phenomena) and obstacles to model interpretation (e.g., systems that lack ground truth constitutive models). Stable attracting sets provide critical insight into the mechanisms underlying the dynamics of biological systems including locomotion [29], gene regulatory networks [45], and neural population dynamics [56]. We thus expect broad utility of VERT for learning models of biological and engineering systems, where the behavior of observed trajectories is strongly influenced by persistent disturbances and internal closed loop control processes.
@@
We demonstrated the efficacy of VERT for accurately identifying attracting sets in synthetic data from dynamical systems with both parallel (Fig. 3) and sequential (Fig. 4) hierarchies, validated its ability to find posture submanifolds carrying previously discovered template dynamics from experimentally measured human running kinematics (Fig. 5). Our results show that when subsystems with different characteristic time scales are simultaneously active, VERT estimates the distance to the attracting set of the locally dominant subsystem, affording sensitivity to and potentially inference of such parallel hierarchies. In systems with hybrid or jump-like behaviors, VERT detects the locally stable attracting sets on continuous subdomains of the global attracting set. In the human running data, which exhibits features of both parallel and sequential dynamical hierarchy, this sensitivity suggests inference of potential control modules during the stance phase of running. The sublevel sets of the distance estimator accurately predict the phases of stance where ground truth kinematic observables exhibit template behavior, e.g., SLIP-like leg length dynamics and vertical GRF profile. Mounting experimental evidence suggests sensitivity of muscle spindle primary afferents to the rate of change of loading force as a mechanism for reflexive stabilization of postural perturbations [36]. The VERT predictions of post-transient template phases of stance are characterized by linear or constant accelerations (i.e., zero yank) of the ’template’ kinematic variables for the respective arm-torso and leg control modules. These observations support the possibility of using VERT to accurately predict reflexive muscle activation dynamics from motion capture data, a prospect with potential breakthrough applications for biomechanics, robotics, and simulation of human movement. The identification of distinct transient timescales further suggests that VERT can be used in experimental design to develop perturbation assays that untangle the roles of distributed postural control modules for stabilizing locomotion. The finite time Lyapunov exponent has served as the standard metric of dynamic stability in both human [18] and animal [1] locomotion. The fiberwise distance estimate afforded by VERT is fundamentally distinct from the information contained in the Lyapunov spectrum (see SI Sec. 2 for further discussion). We expect that VERT will allow researchers to take the next step towards probing the specific mechanisms that give rise to the global stability of observed locomotion behaviors.
@@
Integration of the VERT pipeline with tools for geometric and topological inference is a natural next step — discovered posture submanifolds likely (e.g., almost always in robotics) have dimensions greater than $1$, and are almost always highly nonlinear. This motivates developing a funnel to the computational pipes of Topological Data Analysis [5], allowing one to recognize, in a model-agnostic way, the topological structure of the underlying attracting set. As the attracting sets are still expected to be of relatively low dimensions, we conjecture that the available topological inferences (primarily, ranks of the homology groups) would be sufficient to achieve this goal. Once the underlying attracting set is recovered, the estimation of the template behavior — the dynamics on the attractor — could be achieved by the standard approximation tools [57, 61]. The promising results from the A-HM and A-vdP models motivate further work on inferences for systems with multiscale behavior. We showed that applying VERT to hypothesis-informed projections of the human postural state can provide insight into the dynamics of particular appendages and body segments. Learning such projections directly from the outputs of VERT is an exciting prospect. To ensure our analysis of the human running data was refined to the stance phase, we used the independently measured vertical GRF signal to segment the full trial into phases when the right leg was in contact. In principle, integrating detection of hybrid reset events into VERT should be possible. The saltation matrix characterizes the first order variational update at a reset [32], and can be estimated from the local sample covariance matrices, which are already computed and stored by VERT to estimate the vielbein. Future work will explore the prospect of using such estimates to localize guard surfaces in state space.
@@
Trajectories for the A-HM model were produced by integrating the model defined by Eqns. 2 & 3 using MATLAB’s ode45 function for the integration interval $[0:0.05:10]$. 10 polar coordinate pairs were sampled from the continuous uniform distributions $\rho\sim\mathcal{U}(2,4)$ and $\theta\sim\mathcal{U}(0,2\pi)$. For each polar coordinate pair $(\rho,\theta)$, five initial conditions were computed by sampling: $(\rho,\theta)+\mathcal{N}(0,0.3)$, producing a dataset of 50 trajectories. Trajectories were transformed to Cartesian coordinates before passing to VERT. To produce the plots in Fig. 3, control parameter combinations $(\gamma_{\rho},\gamma_{\delta})$ were produced from a uniformly spaced $50\times 50$ grid with $\gamma_{\rho},\gamma_{\delta}\in[1,3]$. VERT was used to compute the distance estimator for 10 trajectories. The MSE residual was computed over these 10 trajectories. For each control parameter combination, we performed 5 trials each with independently sampled initial conditions, and the MSE residuals for each trial were averaged to produce Figs. 3C,D. Details for computing the MSE residuals are described in Section 5D of the supplement. Parameters used to create the plots in Fig. 3 are available in Section 5C of the supplement.
@@
Trajectories for the A-VDP model were produced by integrating the model defined by Eqn. 4 using MATLAB’s ode45 function for the integration interval $[0:0.01:20]$. To produce the plots in Fig. 4, 20 initial conditions were independently sampled from the normal distributions: $x_{1}\sim\mathcal{N}(0.5,0.2)$, $x_{2}\sim\mathcal{N}(0.3,0.2)$, $x_{3}\sim\mathcal{N}(0.5,0.2)$, $x_{4}\sim\mathcal{N}(0.1,0.2)$. The numerical derivative $\dot{x}_{2}$ plotted in Fig. 4 was obtained using first order finite differences.
@@
[…]